OpenStudy (anonymous):

Is the line with equation x = 4 horizontal or vertical? What is the slope of this line?

OpenStudy (jdoe0001):

x = 4 means \( \begin{array}{llll} y = 10 & x = 4\\ y = 100 & x = 4\\ y = -10 & x = 4\\ y = -100 & x = 4\\ y = 1,000,000 & x = 4\\ y = \pm \infty& x= 4 \end{array}\) so 

OpenStudy (jdoe0001):

what's the slope for this line, well, slope is rise/run run will be \(\bf x_2-x_1\) but notice, that x is always the same value, in this case 4 so the run is \(\bf 4-4 = 0\), that makes our denominator in the equation \(\bf \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \qquad \color{red}{undefined}\)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."